(k-1)x+3k=ax+24
If the equation above is true for all values of x, where k and a are constants, what is the value of a ?

Answer by ikleyn(52780)

.
It is the coefficient at "x" in the left side: a = k-1.

And 3k = 24, i.e. k = 8.

Hence, a = k-1 = 8-1 = 7.